OpenStudy (anonymous):

Write the equation of the circle whose center is at the origin and whose diameter is 12.

OpenStudy (tkhunny):

Well, do it! What's your best guess?

OpenStudy (anonymous):

D = Pi R(sqd) 12 = Pi (6^2) 12 = Pi (36) 12/36 = Pi 1/3 = Pi

OpenStudy (tkhunny):

?? @nfgallimore What? Diameter is 12. What is the Radius?

OpenStudy (anonymous):

diameter is 12, so radius would be 6 right? So maybe x^2+y^2=blank, I know it is less than 4 but more than three I think.

OpenStudy (tkhunny):

You have it. "blank" is \(r^{2} = 36\)

OpenStudy (anonymous):

so how would this formula work out though? Would it come out like x^2+y^2=36?

OpenStudy (tkhunny):

That's it!! Why do you doubt?

OpenStudy (anonymous):

haha, because I'm not the best at geometry xD thank you!!

OpenStudy (tkhunny):

Well, there you go. Now you're a little bit better. Hang in there!